You must estimate the mean temperature (in degrees Fahrenheit) with the following sample temperatures:

60.4
80.3
63.9
49.5
77.5
66.3



Find the 90% confidence interval. Enter your answer as an open-interval (i.e., parentheses) accurate to two decimal places (because the sample data are reported accurate to one decimal place).

90% C.I. =

Answer should be obtained without any preliminary rounding.
Answer #1

Answer: Given data X = 60.4,80:3, 63:9,49.5, 11:5, 66:3 N=9 We have to find 90% confidence interwal estimate for population m0.1 Here CL = 0.90 . CL=1-a a = 1-CL => 1-0.90 critical value i ta, df = t o4,5 =62.015 76 V6 .: P[66-317 -2015 X 351 <H <66